# Entity or Relation Embeddings? An Analysis of Encoding Strategies for Relation Extraction

### v1

@@ 4.3 Analysis
TACRED F1 74.2 74.7 77.6 77.9
TACREV F1 74.8 76.8 82.9 83.4
Re-TACRED F1 86.0 87.4 91.5 91.9
Wiki-WD F1 79.2 84.3 89.1 89.4
NYT-FB P@10 93.7 93.9 94.2 94.5
NYT-FB P@30 93.0 92.3 92.9 93.3
@@
Table 4: Comparison of different language models, using the [H,T]+Mask relation embedding strategy under standard pre-training.
Sentence Label
Asia Bibi was sentenced to hang in Pakistan’s central province of Punjab earlier this month after being accused of insulting the Prophet Mohammed in 2009.
Gold: no relation
[H,T]: per:state or provinces of residence
[H,T]+Mask: no relation
“To get rich is glorious" has been the mantra in booming communist China for 30 years, but few have embraced the slogan more vigorously than Wen Qiang, a leading law official in the country’s southwest.
Gold: no relation
[H,T]: per:countries of residence
[H,T]+Mask: no relation
Benjamin Chertoff is the Editor in Chief of Popular Mechanics magazine as well as the cousin of the Director of Homeland Security, Michael Chertoff.
Gold: no relation
[H,T]: per:employee of
[H,T]+Mask: no relation
WASHINGTON – The National Restaurant Association gave \$35,000 – a year’s salary – in severance pay to a female staff member in the late 1990s after an encounter with Herman Cain, its chief executive at the time, made her uncomfortable working there, three people with direct knowledge of the payment said on Tuesday.
Gold: org:top members/employees
[H,T]: no relation
[H,T]+Mask: org:top members/employees
Table 5: Comparison of the TACREV test set outputs from the [H, T] and [H, T] + Mask models that were initialised using RoBERTa-large.
Comparison of Different Language Models
In Table 4, we compare the performance of roberta-large, which we have been using for our main experiments, with three other LMs: bert-base-uncased [^5] https://huggingface.co/docs/transformers/model_doc/bert, bert-large-uncased and albert-xxlarge-v1666https://huggingface.co/docs/transformers/model_doc/albert. Unsurprisingly, we can see that bert-large-uncased outperforms bert-base-uncased. We furthermore see, across all datasets, that roberta-large outperforms bert-large-uncased, but the best results are obtained by albert-xxlarge-v1. This is consistent with the findings from Zhong and Chen (2021). The main advantage of albert-xxlarge-v1 is that it uses 4096-dimensional embeddings, compared to 1024-dimensional embeddings for bert-large-uncased and roberta-large, despite being smaller than the latter two models (due to the use of a parameter sharing across layers). This allows the model to learn more informative relation embeddings.
Qualitative Analysis of Model Outputs
In Table 5, we compare predictions of the [H,T] and [H,T]+Mask models, for examples from the TACREV test set. These examples illustrate how the [H,T] model often overly relies on the semantic types of the entities, without fully taking into account the actual sentence context. As the first two examples illustrate, given a person and a country, the [H,T] model often predicts that the person resides in that country, even if there is no relationship. The third example shows a similar issue, which arises when the sentence refers to a person and an organisation, in which case the [H,T] model often incorrectly predicts that the person is employed by that organisation. As the fourth example illustrates, however, the opposite situation also arises, where [H,T] mistakenly predicts that there is no relation. These examples support the view that the [H,T] model focuses almost exclusively on modelling the entity types, which is not always sufficient. To further illustrate this point, we present an analysis of the confusion matrices in the appendix.
5 Conclusions
Limitations
Our analysis in this paper was limited to the English language, and we only considered the setting of fully supervised sentence-level relation extraction. Since our focus was on learning representations (i.e. relation embeddings), it is not straightforward to transfer our findings to problems such as zero-shot or few-shot relation extraction. We did not consider the use of LLMs for relation extraction in this paper. On the one hand, this is due to the fact that applying LLMs to this setting is not straightforward, especially when the set of candidate relation labels is large. Progress in this area has indeed been slow, as we highlighted in the related work section. Moreover, while LLMs can be used to extract symbolic representations (e.g. knowledge graph triples), they are less suitable for learning embeddings. Traditionally, relation embeddings have primarily been used as an intermediate representations, before relation labels were predicted, and from this perspective, we may wonder whether relation embeddings are still needed in the LLM era. However, beyond acting as an intermediate representation, embeddings have a number of important advantages. They can, in principle, capture much more subtle distinctions than is possible with pre-defined discrete relation labels. As such, they are more suitable for modelling relational similarity (i.e. analogy), for instance.
Appendix A Additional Analysis
Comparison of Confusion Matrices
In Figures 1 and 2, we show the confusion matrix for 5 randomly sampled relations and the “no relation” label, for the TACREV test set. For this analysis, we focus on the [H,T] and [H,T]+Mask models. We group any predictions of labels outside the sampled six classes as the “other” class. We can see that the “no relation” label is the most challenging label for both models: most errors arise when the models predict a relationship when there is none, or vice versa. We can also see, however, that the [H,T]+Mask consistently improves on [H,T] when it comes to modelling the “no relation” label. This supports the conclusions from our qualitative analysis.

### v2

@@ Appendix C Ablations and Additional Experiments
Mask TACRED F1 20.9 22.2 23.3 25.1
TACREV F1 20.6 21.7 22.9 24.8
Re-TACRED F1 21.3 22.1 23.2 25.4
Wiki-WD F1 44.9 46.2 48.9 51.3
NYT-FB P@10 48.2 49.5 52.3 53.7
NYT-FB P@30 47.6 48.2 49.7 51.2
[H,T] TACRED F1 58.6 59.3 61.4 62.9
TACREV F1 59.1 60.7 63.6 64.3
Re-TACRED F1 67.4 69.1 72.3 73.5
Wiki-WD F1 72.4 73.1 75.2 76.9
NYT-FB P@10 75.2 76.7 79.6 80.7
NYT-FB P@30 73.9 75.2 78.5 79.3
[H,T,Mask] TACRED F1 70.1 71.6 73.0 74.7
TACREV F1 70.8 71.1 73.8 75.2
Re-TACRED F1 76.3 78.4 81.9 84.7
Wiki-WD F1 76.2 78.7 81.7 82.4
NYT-FB P@10 80.5 84.3 88.4 90.2
NYT-FB P@30 79.9 82.7 86.1 89.4
[H,T] + Mask TACRED F1 74.8 75.1 78.5 78.9
TACREV F1 75.6 77.3 84.6 83.9
Re-TACRED F1 86.4 87.9 91.9 92.3
Wiki-WD F1 79.7 84.9 89.6 89.8
NYT-FB P@10 93.9 94.1 94.7 94.9
NYT-FB P@30 93.4 92.5 93.1 93.8
@@
Table 10: Comparison of different language models, using different relation embedding strategies under standard pre-training.
F1
1% 5% 10% 100%
Baselines
SpanBERT 0.0 28.8 1.6 70.8
RoBERTa 7.7 41.8 55.1 71.3
LUKE 17.0 51.6 60.6 72.0
$\text{NLI}_{\text{RoBERTa}}^{\dagger}$ 56.1 64.1 67.8 71.0
$\text{NLI}_{\text{DeBERTa}}^{\dagger}$ 63.7 69.0 67.9 73.9
$\text{SURE}_{\text{PEGASUS}}^{*}$ 52.0 64.9 70.7 75.1
Standard pre-training
Mask 8.3 11.6 15.0 23.3
[H, T] 19.7 27.5 33.9 61.4
[H, T, Mask] 48.4 54.6 61.3 73.0
[H, T] + Mask 53.0 65.9 71.7 78.5
EnCore + Mask 52.9 65.8 71.6 78.9
Gigaword pre-training
Mask 9.9 13.1 18.3 24.6
[H, T] 21.1 28.7 34.3 63.2
[H, T, Mask] 49.2 55.1 62.1 74.4
[H, T] + Mask 53.5 66.1 71.9 78.5
EnCore + Mask 53.4 66.2 72.1 79.1
[…]